A rectangle with diagonals of length 20 cm has sides in the ratio 2 : 1. Find the perimeter
Answer by checkley77(12844)

X^2*2X^2=20^2
2X^4=400
X^4=400/2
X^4=200
X=FOURTHRT200
X=3.76 ANS. FOR THE SHORT SIDE
2*3.76
7.52 ANS. FOR THE LONGER SIDE.
PROOF:
3.76^2*2*3.76^2=20^2
14.1376*2*14.1376=400
14.1376*28.2752=400
400~400